# ]organization=LASIGE, Faculdade de Ciências, Universidade de Lisboa , country=Portugal

### 26 Jan 2024 v1

@@ 7.1.2 Execution time
In our findings, the execution times of the models varied significantly. The implementation process involves running LLMs on a GPU server, which requires several days of continuous execution. GPT4all took the longest, with a total execution time of 132 hours and five minutes. Falcon followed with 54 hours and 3 minutes. Stanford Alpaca completed its tasks in 15 hours and 48 minutes, while Dolly had a slightly shorter execution time of 13 hours and 38 minutes. ChatGPT-4 was more efficient, ranking third with an execution time of 11 hours and 50 minutes. The fastest among them was Alpaca-LoRA, with an execution time of 10 hours and 12 minutes. The models with versions of 30B and 65B parameters exhibit longer execution times when answering questions. An impressive aspect of the Alpaca-LoRA model with 65B parameters is that despite its significantly larger parameter count than Dolly, which has 7B and 12B parameters, both models exhibit equal execution times and F1-score.
The last row of Table 1 is devoted to the multi-task model [5], which we discuss in the related work section. This model is a Bidirectional Long Short-Term Memory (BiLSTM) trained for binary classification tasks, as well as NER. It achieved an accuracy of 0.94, which is equivalent to the accuracy of the ChatGPT model. However, the execution time is 43 minutes, which is much shorter than that of the chatbot models. We have highlighted this trained model for specific tasks due to its use of the same Twitter dataset introduced for training. It is included here as a baseline. Based on the experimental results, it is evident that ChatGPT with a context window size of 512 outperforms the open-source models in terms of the F1-score. One key contributing factor to this accuracy gap is the significantly larger context window size employed by the ChatGPT.
@@ 7.2 Evaluation of NER
Table 2 shows the results for identifying the ‘product version’ entity, which is 0.54 of F1-score, based on 2550 requested questions. This F1-score exceeds the corresponding value for identifying the ‘organization name’ entity. Our analysis for the ESP approach is based on the ChatGPT version released on July 13th, 2023. The results in the last line of Table 2 indicate that the calculated F1-score for 11074 questions is unexpectedly low, reaching 0.10 accuracy. The findings demonstrate a significant deviation from the current NER outcomes as presented in the [5] study, which achieved a 0.94 F1-score. Our GLP approach evaluation is based on the ChatGPT released on August 2, 2023.
@@
Table 2: Comparison of NER Task accuracy Achieved by the ChatGPT-4 (8k context) Model Using Two Different Approaches
@@
F1-score Execution Time
ESP 2550
@@
8 Challenges and Limitations
Chatbots present opportunities in numerous applications but also face challenges and limitations that hinder their effective utilization. This section explores the challenges and limitations of binary classification and NER tasks.
8.1 Binary Classification
In Section V.C., we discuss how each chatbot model displayed unique response behaviors when answering questions. This necessitated a cleaning step after collecting responses to ensure answer consistency, particularly because our goal was to obtain a precise binary (yes or no) response. However, some responses implied a ‘no’ or a ‘yes’ without explicitly using these words. For instance, a model might answer the question, ‘This is not related to cybersecurity’. Consequently, it is crucial to review and validate the answers in the output file. Our evaluation process involved two key steps to ensure reliability and accuracy. First, we implemented an automated validation method for each output file to confirm the presence of ‘yes’ or ‘no’ responses at the beginning of the responses. This helped to filter out potentially incorrect answers, such as those lacking an explicit ‘no’ or ‘yes’. Second, we conducted a thorough manual review of each response to confirm its accuracy and alignment with the expected answers. In cases such as the aforementioned example, we manually annotated the response with ‘no’ to correctly classify it. This comprehensive manual validation adds a crucial layer of scrutiny, bolstering the accuracy and reliability of the results. While this comprehensive validation process enhances the accuracy of our binary classification, it also introduces another challenge: the need for timely response processing. Timeliness is a critical factor in the use of chatbot models for binary classification tasks, particularly in CTI applications. The demand for real-time processing is essential, as any delay in classifying responses can significantly impact decision-making.
8.2 Named Entity Recognition
@@ 9 Discussion
Although our evaluation primarily focused on a specific Twitter dataset, it is worth noting that numerous other OSINT resources, such as LinkedIn and Facebook, remain unexplored. Moreover, it is important to note that a manual review of chatbot responses that follow the automated checking phase may introduce potential human errors into the assessment process. In addition, it is time-consuming and thus not feasible for day-to-day operation.
@@ 10 Conclusion
Our study highlights the capability of LLM-based chatbots to recognize cybersecurity-related tweets from a Twitter dataset. Although they sometimes operate without direct training data, their adaptability to new and unfamiliar contexts stands out. Both the open-source and commercial chatbot models performed similarly to the trained models in identifying security-related texts, often achieving the same level of performance.
To conclude our research on NER tasks, several key findings arose from our experiments with chatbot models. First, the ability of current chatbots to detect cybersecurity entities still has limitations. Even when training on vast datasets, these models did not yield the desired performance on our specific test data. This highlights the need for further focused research and refinement in the precise application of chatbot models to identify cybersecurity entities, despite their significant advancements in various fields. Although ChatGPT is not as accurate as fine-tuned NER models and has significant flaws, it can be used as a primary tool for text pre-annotation.

### 13 Mar 2024 v2

@@ 6.1.2 Execution time
In our findings, the execution times of the models varied significantly. The implementation process involves running LLMs on a GPU server, which requires several days of continuous execution. GPT4all took the longest, with a total execution time of 132 hours and five minutes. Falcon followed with 54 hours and 3 minutes. Stanford Alpaca completed its tasks in 15 hours and 48 minutes, while Dolly had a slightly shorter execution time of 13 hours and 38 minutes. ChatGPT-4 was more efficient, ranking third with an execution time of 11 hours and 50 minutes. The fastest among them was Alpaca-LoRA, with an execution time of 10 hours and 12 minutes. The models with versions of 30B and 65B parameters exhibit longer execution times when answering questions. An impressive aspect of the Alpaca-LoRA model with 65B parameters is that despite its significantly larger parameter count than Dolly, which has 7B and 12B parameters, both models exhibit equal execution times and $\text{F}_{1}$ score.
The last row of Table 1 is devoted to the multitask model [41] discussed in Section 2.4 of the related work. Since this model achieves the best performance in OSINT-based CTI extraction, it was selected as the reference specialized model to evaluate the chatbots. Furthermore, the Twitter dataset employed for both this model and the chatbots is identical, guaranteeing a fair evaluation under comparable conditions.
This model is a Bidirectional Long Short-Term Memory (BiLSTM) trained for binary classification tasks and NER. It achieved an accuracy of 0.94, equivalent to the ChatGPT model’s accuracy. However, the execution time is only 43 minutes, which is much shorter than the 11+ hours required by ChatGPT. Based on the experimental results, it is evident that ChatGPT, with a context window size of 512, outperforms the open-source models in terms of the $\text{F}_{1}$ score. One key contributing factor to this accuracy gap is the significantly larger context window size employed by the ChatGPT.
@@ 6.2 Evaluation of NER
Table 2 shows the results for identifying the ‘product version’ entity, which is 0.43 of the $\text{F}_{1}$ score, based on 11074 requested questions. This $\text{F}_{1}$ score exceeds the corresponding value for identifying the ‘organization name’ entity. Our analysis for the ESP approach is based on the ChatGPT version released on July 13th, 2023. The results in the last line of Table 2 indicate that the calculated $\text{F}_{1}$ score for 11074 questions is unexpectedly low, reaching 0.10 accuracy. The findings demonstrate a significant deviation from the current NER outcomes as presented in the [41] study, which achieved an $\text{F}_{1}$ score of 0.94. Our GLP approach evaluation is based on the ChatGPT released on August 2, 2023.
@@
Table 2: Comparison of NER task accuracy achieved by the ChatGPT-4 (8k context) model using two different approaches.
@@
$\text{F}_{1}$ score Execution Time
ESP 11074
@@
6.3 Challenges and limitations
Chatbots present opportunities in numerous applications but also face challenges and limitations that hinder their effective utilization. In this study, we outline several challenges and limitations encountered during our experiments, which are detailed in this section. Additionally, we faced specific limitations when applying chatbots to the specialized CTI tasks, leading to issues with timeliness that are elaborated upon in Section 6.3.1. A common challenge to binary classification and NER was the generation of effective prompts, which required progressive refinement until satisfactory performance was achieved. This section explores the challenges and limitations encountered in the binary classification and NER tasks.
6.3.1 Binary classification
In Section 4.2.1, we discuss how each chatbot model displayed unique response behaviors when answering questions. This necessitated a cleaning step after collecting responses to ensure answer consistency, particularly because our goal was to obtain a precise binary (yes or no) response. However, some responses implied a ‘no’ or a ‘yes’ without explicitly using these words. For instance, a model might answer the question, ‘This is not related to cybersecurity’. Consequently, it is crucial to review and validate the answers in the output file. Our evaluation process involved two key steps to ensure reliability and accuracy. First, we implemented an automated validation method for each output file to confirm the presence of ‘yes’ or ‘no’ responses at the beginning of the responses. This helped to filter out potentially incorrect answers, such as those lacking an explicit ‘no’ or ‘yes’. Second, we conducted a thorough manual review of each response to confirm its accuracy and alignment with the expected answers. In cases such as the aforementioned example, we manually annotated the response with ‘no’ to correctly classify it. This comprehensive manual validation adds a crucial layer of scrutiny, bolstering the accuracy and reliability of the results. While this comprehensive validation process enhances the accuracy of our binary classification, it also introduces another challenge: the need for timely response processing. Timeliness is a critical factor in the use of chatbot models for binary classification tasks, particularly in CTI applications. The demand for real-time processing is essential, as any delay in classifying responses can significantly impact decision-making.
6.3.2 Named entity recognition
@@ 6.4 Discussion
Although our evaluation primarily focused on a specific Twitter dataset, it is worth noting that numerous other OSINT resources, such as blog posts and security forums (even in the dark web), remain unexplored. Moreover, it is important to note that a manual review of chatbot responses that follow the automated checking phase may introduce potential human errors into the assessment process. Further, such verification is time-consuming and thus not feasible in the day-to-day operation of a security operating centre.
@@ 7 Conclusion
We assess the capabilities of open-source and paid LLM-based chatbots to recognize cybersecurity-related tweets and extract pertinent information from these. Both types of models can perform similarly to specialized models trained specifically for the binary classification task of identifying cybersecurity-related tweets, often achieving the same level of performance. On the contrary, the chatbot models’ performance is still very poor on named entity recognition to extract security elements from tweets. Even when training on vast datasets, these models did not perform comparably to specialized models on the test data.
Our results highlight the need for further research and refinement in the application of LLM chatbot models to extract threat indicators from open-source intelligence. Although open-source and paid chatbots compared evenly with specialized trained models on cybersecurity binary text classification, they fell below acceptable performance on named entity recognition. Moreover, they can not compete with specialized models on timeliness and cost.
Based on our study, we identify various possibilities for future work based on the following research questions:
  1. 1.
    How can LLM chatbots be further optimized for cost-effective real-time CTI detection on social media platforms?
  2. 2.
    How to improve the NER capability of chatbots for the extraction of indicators of compromise?
  3. 3.
    How can cybersecurity specialists’ feedback be used to increase the efficiency and cost-effectiveness of open-source chatbots?
@@ Acknowledgement
Appendix Appendix A
The template for the guideline prompt used in the ChatGPT GLP NER approach is shown in Fig A1.
Fig. A1: GLP NER approach: employing a ChatGPT-4 guideline prompt template
Appendix Appendix B
As described in Section 4.1, for each collected tweet a dataset entry is generated, including timestamp, keywords, original tweet, pre-processed tweet, cybersecurity relevance binary label, and sequence of named entities in the pre-processed tweet. Table B1 presents two examples of dataset entries. In the relevance column, ‘1’ denotes an entry considered relevant for cybersecurity, a ‘0’ means otherwise. The last column shows the tags used to label the different NER entities.
Table B1: Samples from the 31281 tweet entries in the dataset
timestamp keywords
original tweet
pre-processed tweet relevance entities
2018-07-24 01:00:46+00:00 oracle
RT Oracle: Learn to use and understand #Oracle’s Internet Intelligence Map https://t.co/l06Nyf1FFF Dyn https://t.co/uzozFKwm97
rt oracle learn to use and understand oracle s internet intelligence map dyn 0 -
2016-12-09 19:19:38+00:00 internet explorer
threatmeter: [dos] - Microsoft Internet Explorer 9 MSHTML - CDisp Node::Insert Sibling Node Use-After-Free (MS13-0… https://t.co/gLvEwpDL9v
threatmeter dos microsoft internet explorer 9 mshtml cdisp node::insert sibling node use-after-free ms13-0 1 O O B-ORG B-PRO I-PRO B-VER O O O O O B-VUL B-ID
Appendix Appendix C
The confusion matrices are provided the test and model combinations considered, excluding the 7B parameter models in Table 1 which achieved the worse results in the respective group. The rows in the matrices correspond to the actual expected result, whereas the columns show the predicted results.
Fig. C1: Confusion matrix of binary classification task